OpenStudy (anonymous):

find the domain of this function? explain please. noncalculator f(t)= ∛(t-1)

OpenStudy (zzr0ck3r):

we are looking for what the x's can be. is there any reason that x cant be a certain number or numbers?

OpenStudy (anonymous):

i honestly don't know

OpenStudy (anonymous):

it can be anything so its just negative infinity and infinity

OpenStudy (anonymous):

(-oo,oo)

OpenStudy (zzr0ck3r):

yes (-infinity,infinity)

OpenStudy (zzr0ck3r):

or {x:x is an element of R}

OpenStudy (anonymous):

thanks sometimes these are so easy i over think them haha

OpenStudy (zzr0ck3r):

np, the only time you will run into problems with roots is when the root is even, do you understand why?

OpenStudy (anonymous):

no could you explain

OpenStudy (anonymous):

like there is one ezample

OpenStudy (anonymous):

quadroot of (x^2-6x)

OpenStudy (anonymous):

i think its called a quad root hahaha

OpenStudy (zzr0ck3r):

right, so what is the domain of \[\sqrt[4]{x}\]?

OpenStudy (kainui):

You can't take the square root of a negative number because you can't ever multiply something by itself and get a negative number! -1*-1=1 and 1*1=1 You also can't divide by 0.

OpenStudy (zzr0ck3r):

so notice sqrt(-1) = 4th root of(-1) = 6th root of (-1) = i

OpenStudy (anonymous):

couldnt it be -oo,oo

OpenStudy (zzr0ck3r):

no, can you take the 4th root of -5?

OpenStudy (zzr0ck3r):

in other words a*a*a*a = -5 whats a?

OpenStudy (zzr0ck3r):

\[\sqrt[even number]{x} \] is undefined when x <0 \[\sqrt[odd number]{x} \] is defined for all real numbers (notice i is not a real number)

OpenStudy (precal):

 look at the graph, notice all of the x's are used on the number line that is why it is "all real numbers" for the domain aka as (-infinity, +infinity)

OpenStudy (anonymous):

i'm still lost.

OpenStudy (anonymous):

sorry

OpenStudy (precal):

 ok lets do one this is a line segment